Why is work done negative only when cos theta is 90<theta<= 180 but not when it is 90<theta <270 .
1
Expert's answer
2020-11-17T11:25:52-0500

Answer

Work done by force is given by

"W=Fdcos\theta"

Work will be negative or positive it will depend upon the value of "cos\theta"

When "90°<\theta<270°" then value of cos"\theta" will be negative so work done will be negative.

When value of theta "0°<\theta<90° and \space 270°<\theta<360°"

Then value of cos"\theta" will be positive.

So work done will be positive.